OpenStudy (anonymous):

Match each of the following with its correct simplified solution. ab(a + b) (a + b)(2a + 2b) (a + b)2 (a + b)(a - b) match. A. a2 + 2ab + b2 B. 2a2 + 4ab + 2b2 C. a2 - b2 D. a2b + ab2 @timo86m @primeralph @Directrix

OpenStudy (anonymous):

ab(a + b) ba^2+ba^2 can you do more to it? to make it match one of the 3 choices?

OpenStudy (anonymous):

(a+b)^2=a^2+2ab+b^2 (a+b)(a-b)=a^2_b^2 ab(a+B)=a^2b+ab^2 (a+b)(2a+2B)=2a^2+4ab+2b^2

OpenStudy (anonymous):

how is that @ green bean

OpenStudy (anonymous):

thank you so much, that's great you have no idea how much your helping me

OpenStudy (anonymous):

sorry i messed up. and then OS broke on me lol.